How do I solve the following equation for "T"?
F=1/T

Answer by ptaylor(2198)

F=1/T Lets start by multiplying both sides of the equation by T
F(T)=(1/T)(T)
FT=1 Next divide both sides by F
FT/F=1/F or
T=1/F----ans

Hope this helps ----ptaylor